OpenStudy (anonymous):

What is the r value of the following geometric sequence? -81, 54, -36, 24,…

OpenStudy (freckles):

common ration can be found by doing term divided by previous term

OpenStudy (freckles):

54/(-81)=? -36/(54)=? 24/(-36)=? should all be the same if this is a geometric sequence

OpenStudy (anonymous):

-2/3?

OpenStudy (zzr0ck3r):

correct